17. Below is the distribution of violent crimes in New York city in 2015 and the...

17.

Below is the distribution of violent crimes in New York city in 2015 and the breakdown of a sample of 1000 violent crimes in New York city in this year. Assume 5% level of significance.

Violent crime Count in the study Percent in 2015
Murder 5   1%
Rape 50 4%
Robbery 352 35%
Assault 593 60%
Total 1000 100%


Under the null hypothesis, what would be the expected breakdown for each type of violent crime (murder, rape, robbery, assault) be?

Question 17 options:

A)

1; 4; 35; 60

B)

10; 40; 350; 600

C)

0.05; 2; 123.2; 355.8

D)

250; 250; 250; 250

E)

5; 50; 352; 593

18.

In a study of cereal leaf beetle damage in oats, researcher measured the number of beetle larvae per stem in small plots of oats after randomly applying one of two treatments: no pesticide or malathion at the rate of 0.25 lb. per acre. Here are the summary statistics:

Group Treatment n s
1 mg Control 13 3.47 1.21
2 mg Malathion 14 1.36 0.52



To construct a 95% confidence interval for the difference in mean numbers of beetle larvae per stem for the two treatments, you need to calculate the standard error. Which of the following values provides the standard error?

Question 18 options:

A)

-0.363

B)

0.363

C)

0.130

D)

0.305

E)

0.383

19.

Core Plus Mathematical Project is a new approach to teach Mathematics. A teacher was interested in finding whether the performance of students who were taught using CPMP is better than the performance of students who were taught using the traditional curriculum. Scores of 300 CPMP students were compared with the scores of 350 students in a traditional Math program. The teacher was interested in seeing if there was a difference in mean scores of the two groups. What type of test should they perform?

Question 19 options:

A)

Two sample independent test for the difference in proportions

B)

One sample independent test for the difference in a proportion

C)

Two sample independent test for the difference in means

D)

One sample independent test for the difference in a mean

E)

Two sample dependent test for the difference in means
Answer #1

17.
expected breakdown for each type of violent crime = E = npi
= 1000 * 1%, 1000 * 4%, 1000 * 35%, 1000 * 60%
= 10, 40, 350, 600

B)

10; 40; 350; 600


18.

Standard error = \sqrt{ s_1^2 /n_1 + s_2^2 /n_2} = \sqrt{1.21^2/13 + 0.52^2/14} = 0.363

B)

0.363

19.

There are two independent samples - students who were taught using CPMP and students who were taught using the traditional curriculum
We need to test for a difference in mean scores of the two groups.

So, the appropriate test is,

C)

Two sample independent test for the difference in means